OpenStudy (anonymous):

Multiply the following fractions and write your answer in simplest form. 12 3/5 x 1 3/7. A. 20 B. 630/35 C. 24 D. none of these

OpenStudy (anonymous):

I thought it was B.

OpenStudy (anonymous):

its not B because its not in simplest form.

OpenStudy (anonymous):

oh okay sorry

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

\[\left(12+\frac{3}{5}\right) \left(1+\frac{3}{7}\right)=\frac{63\ 10}{5\ 7}=18 \]

OpenStudy (anonymous):

its D. none of these

OpenStudy (anonymous):

18 isn't one of my answers though so the answer is D?

OpenStudy (anonymous):

yep. B is almost it. when its simplified and rounded to the nearest whole number, it would be 18

OpenStudy (anonymous):

but its not addition its multiplication

OpenStudy (anonymous):

when you have a whole number and a fraction, you have ton add them together before you multiply, which is what he did^

OpenStudy (anonymous):

oh okay thanks you 2 :)